### v1

# Two classes of oligomorphic groups with smooth topological isomorphism relation

Abstract.
It is an open question whether topological isomorphism of oligomorphic groups is smooth in the sense of Borel reducibility. We show smoothness for two Borel classes of oligomorphic groups: groups with no algebraicity, and groups with finitely many essential subgroups up to conjugacy. This second class, which is new, contains the class of automorphism groups of $\aleph_{0}$-categorical structures that have weak elimination of imaginaries, studied in (G. Paolini, the isomorphism problem for oligomorphic groups with weak elimination of imaginaries, Bull. London Math. Soc., 2024). Towards showing this, we obtain results on general oligomorphic groups $G$ that are of independent interest. $\mathrm{Inn}(G)$ is closed in $\mathrm{Aut}(G)$, where $\mathrm{Aut}(G)$ carries the topology of pointwise convergence on open cosets. $N_{G}/G$ is profinite, where $N_{G}$ denotes the normaliser of $G$ in $\mathrm{Sym}(\omega)$. The topological group $\mathrm{Out}(G)=\mathrm{Aut}(G)/\mathrm{Inn}(G)$ is totally disconnected, locally compact. Finally, we show that if $G$ is in one of our two classes, then $\mathrm{Aut}(G)$ is topologically isomorphic to an oligomorphic group, and that $\mathrm{Out}(G)$ is profinite.
The first author was supported by the Marsden fund of New Zealand, grant number UOA-1931. The second author was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA, and by INdAM Project 2024 (Consolidator grant) “Groups, Crystals and Classifications”.
1. Introduction
Borel reducibility was introduced in [10] as a means of comparing the complexity of equivalence relations $E$ and $F$ on Borel spaces $X$ and $Y$, respectively. One says that $E$ is Borel reducible to $F$, written $(X,E)\leqslant_{B}(Y,F)$, if there is a Borel function $f\colon X\to Y$ such that $x_{0}Ex_{1}\leftrightarrow f(x_{0})Ff(x_{1})$ for each $x_{0},x_{1}\in X$. An equivalence relation $E$ on a Borel space $X$ is called smooth if $E$ is Borel reducible to the identity on a standard Borel space such as $\mathbb{R}$. This means that one can in a Borel way assign reals as complete invariants to the equivalence classes of $E$.
In this paper all isomorphism between topological groups will be topological. The set of closed subgroups of $\mathrm{Sym}(\omega)$ forms a Borel subset of the Effros space of closed subsets of $\mathrm{Sym}(\omega)$, and thus can be seen as a standard Borel space (see, for instance, [15, Section 2] for detail). Let $\mathcal{B}$ be a Borel class of closed subgroups of $\mathrm{Sym}(\omega)$ that is invariant under conjugation. We say that $\mathcal{B}$ is smooth if the (topological) isomorphism relation on $\mathcal{B}$ is smooth. For example, the class of topologically finitely generated compact subgroups of $\mathrm{Sym}(\omega)$ is smooth: such groups are given by the collection of isomorphism types of their finite quotients [16], and the set of types of such quotients can be seen as an element of a standard Borel space, when suitably encoded as a set of natural numbers. In contrast, isomorphism on the class of all compact subgroups of $\mathrm{Sym}(\omega)$ is Borel equivalent to the isomorphism relation between countable graphs, and hence not smooth [15, Thm. 4.3]. (Note that the compact closed subgroups of $\mathrm{Sym}(\omega)$ are, up to isomorphism, precisely the countably based profinite groups.)
Recall that a closed subgroup of $\mathrm{Sym}(\omega)$ is called oligomorphic if for each $k<\omega$, its natural action on $\omega^{k}$ only has finitely many orbits. These are precisely the automorphism groups of $\aleph_{0}$-categorical structures with domain $\omega$. The isomorphism relation $\cong_{\mathcal{C}}$ on $\mathcal{C}$ is Borel reducible to a Borel equivalence relation with all equivalence classes countable, and hence far below countable graph isomorphism [18] (see [18, 20] for background). In [18] it was asked whether there is a better upper bound on the complexity of $\cong_{\mathcal{C}}$. Despite attempts by various groups of researchers, so far no better bound has been found. In particular, it remains open whether $\cong_{\mathcal{C}}$ has the lowest possible complexity among the Borel equivalence relation with uncountably many classes:
Question 1.1.
Is the class of oligomorphic groups smooth?
A topological group is called quasi-oligomorphic [18] if it is (topologically) isomorphic to an oligomorphic group. The quasi-oligomorphic closed subgroups of $\mathrm{Sym}(\omega)$ form a Borel class. Its isomorphism relation is Borel equivalent to the isomorphism relation on the class of oligomorphic groups by [18, Section 4.3]. This means that it is no different to answer Question 1.1 for the larger class of quasi-oligomorphic closed subgroups of $\mathrm{Sym}(\omega)$.
Given the apparent difficulty of the problem, it makes sense to restrict it to interesting Borel subclasses of all the oligomorphic groups, and show smoothness for them. Such a pathway was already suggested in [18] for the class of automorphisms groups of $\aleph_{0}$-categorical structures in a finite signature; it was followed in [20] for automorphism groups of $\aleph_{0}$-categorical structures with weak elimination of imaginaries.
1.1. The two smooth classes of oligomorphic groups
We present two classes of oligomorphic groups that are smooth: (A) the groups with no algebraicity, and (B) the groups with finitely many conjugacy classes of essential subgroups. The notion of essential subgroups is introduced in the present work; broadly speaking, these are the open subgroups whose conjugates form a sub-basis of neighbourhoods of the identity, and that given this condition are contained in the fewest possible open subgroups. The class of automorphism groups of $\aleph_{0}$-categorical structures with domain $\omega$ that have weak elimination of imaginiaries is smooth, as proved by the second author in [20]. We will show that this class is contained in the class (B). The class (B) is closed under isomorphism of topological groups, which corresponds to bi-interpretations of the corresponding $\aleph_{0}$-categorical structures. So the larger class of automorphism groups of $\aleph_{0}$-categorical structures with domain $\omega$ that have weak elimination of imaginaries after adding finitely many imaginary sorts is also contained in the class (B), and we obtain a slightly more general result than the one from [20].
By $G$ we will always denote an oligomorphic closed subgroup of $\mathrm{Sym}(\omega)$. For a finite set $S\subseteq\omega$, by $G_{(S)}$ one denotes the pointwise stabiliser of $S$, which is an open subgroup of $G$. If $S$ is a singleton $\{a\}$ we will usually write $G_{a}$ for this.
The class (A) is well-known:
Definition 1.2.
$G$ has no algebraicity if the action of $G_{(S)}$ on $\omega$ has no finite $1$-orbits, for each finite set $S\subseteq\omega$.
Being in this class depends on $G$ as a permutation group. In contrast, the class (B) is determined solely by the lattice of open subgroups, and hence is invariant under topological isomorphism. We need two auxiliary notions, the depth of open subgroups, Def. 1.3, and essential subgroups, Def. 1.4(3).
As is well-known, in the lattice of open subgroups of an oligomorphic group $G$, there are only finitely many elements above any given one; see Fact 2.5 for a proof.
Definition 1.3.
The depth of an open subgroup $U$ is the minimum of all lengths of maximal chains leading from $U$ to $G$. By this definition, $G$ has depth 00.
Definition 1.4.
  1. (1)
    One says that an open subgroup of $G$ is irreducible if it has no proper open subgroup of finite index.
  2. (2)
    We say that an irreducible subgroup $H$ of $G$ is almost essential if every open subgroup of $G$ contains a finite intersection of conjugates of $H$.
  3. (3)
    Such a subgroup $H$ is essential if it is almost essential, and its depth is minimal among the almost essential subgroups.
Our first main result is the following.
Theorem 1.5 (Two smooth classes).
The following Borel classes of oligomorphic groups are smooth:
  1. (A)
    groups with no algebraicity.
  2. (B)
    groups with finitely many essential subgroups up to conjugacy.
We will prove Theorem 1.5 separately for each class: (A) is proved as part of Theorem 4.7(1), and (B) follows from the more general Theorem 5.8(1).
The following notion goes back to Lascar, with a perhaps unfortunate terminology that, however, is well entrenched.
Definition 1.6.
An oligomorphic group is $\mathfrak{G}$-finite if each open subgroup contains a least open subgroup of finite index.
Evans and Hewitt [8] construct canonical examples of oligomorphic groups that are not $\mathfrak{G}$-finite. Bodirsky et al.  [2, Section 3.7] obtained such a group that is the automorphism group of an $\aleph_{0}$ categorical structure in a finite signature; see the last paragraph of that paper for a discussion why the group is not $\mathfrak{G}$-finite. Proposition 5.9 shows that every $\mathfrak{G}$-finite group has an essential subgroup. So the class of $\mathfrak{G}$-finite groups that have only finitely many essential subgroups up to conjugacy forms a Borel subclass of the class in (B), and is therefore smooth. It might be possible a priori that every $\mathfrak{G}$-finite group is of this kind.
As mentioned, Section 5.1 will provide a general result, Thm. 5.8, that implies smoothness of the class (B): Suppose we have a Borel, isomorphism invariant property of open subgroups of oligomorphic groups, denoted by $P$. Then $\mathcal{B}_{P}$ is smooth, where $\mathcal{B}_{P}$ is the class of oligomorphic groups $G$ satisfying the following two conditions:
  1. (i)
    the $P$-subgroups form a sub-basis of neighbourhoods of $1$;
  2. (ii)
    there are finitely many $P$-subgroups up to conjugacy.
To see that 1.5(B) follows, note that the property $P$ of being essential is Borel (as we will verify), and clearly isomorphism invariant. By definition, the conjugates of an essential subgroup form a sub-basis of neighbourhoods of $1$. So the class in (B) is a Borel subclass of $\mathcal{B}_{P}$, and hence smooth as well.
Theorem 1.5 should be seen in the context of an incremental approach towards answering Question 1.1. We do not know whether the isomorphism relation is smooth for the class of all oligomorphic groups, so we search for dividing lines $D$. The ideal situation would be that in the presence of $D$ one has smoothness, and otherwise not; this is analogous to what happens with the dividing lines in classification theory. It would be a strong solution to the whole problem. If a Borel equivalence relation is not smooth, then the relation $E_{0}$ of almost equality between subsets of $\omega$ can be Borel reduced to it by a result of Harrington, Kechris and Louveau (see, e.g., [11, Thm.  6.3.1]). We search for the most general condition ensuring smoothness, so that we can focus on possible constructions for a reduction of $E_{0}$ that exclude groups with that condition.
We think of Theorem 1.5 as providing two dividing lines. The application of (B) we had in mind is for structures with weak elimination of imaginaries after adding finitely many sorts. However, venue (B) and its more general form involving $P$-subgroups are of independent interest, as it is well possible that all $\mathfrak{G}$-finite groups have only finitely many essential subgroups up to conjugacy. All this points in the direction that the “true” dividing line is: $\mathfrak{G}$-finite or without algebraicity. So groups that fail both conditions would be needed to show the class of oligomorphic groups is not smooth. One possible pathway is to consider finite covers of non-$\mathfrak{G}$-finite groups. These little-known groups are considered to be exotic examples of oligomorphic groups. (For an overview on finite covers see [7, Chapter 1].)
1.2. Structural results on $\mathrm{Aut}(G)$ and $\mathrm{Out}(G)$
By $\mathrm{Aut}(G)$ we denote the group of continuous automorphism of $G$, and by $\mathrm{Out}(G)=\mathrm{Aut}(G)/\mathrm{Inn}(G)$ the group of outer (topological) automorphisms of $G$.
Definition 1.7.
Given an oligomorphic group $G$, we define a topology on $\mathrm{Aut}(G)$ by declaring as open the subgroups of the form
$\{\Phi\in\mathrm{Aut}(G)\colon\forall{i=1\ldots n}\ \Phi(A_{i})=A_{i}\},$
where $A_{1},\ldots,A_{n}$ are cosets of open subgroups of $G$.
The coarse group $\mathcal{M}(G)$ of $G$ was introduced in [18]; this is a structure on the domain of open cosets in $G$ with one ternary relation approximating the binary group operation. The duality shown there yields an isomorphism of $\mathrm{Aut}(G)$ and $\mathrm{Aut}(\mathcal{M}(G))$. In particular, the subgroups of $\mathrm{Aut}(G)$ given above form a basis of neighbourhoods of the identity, whence $\mathrm{Aut}(G)$ is a non-Archimedean group. As our second main result, in Theorem 3.3 will show that the normal subgroup $\mathrm{Inn}(G)$ is closed in $\mathrm{Aut}(G)$. Thus $\mathrm{Out}(G)$ is a Polish group as well. The same theorem then shows that $\mathrm{Out}(G)$ has a profinite open subgroup, and hence is totally disconnected, locally compact (t.d.l.c.). For background on such groups see [26, 5].
We now introduce our third main result, Theorem 1.8. Expanding on the methods used in proving Theorem 1.5, for groups in either of the two classes (A) and (B) from 1.5 we will derive the following structural results: $\mathrm{Aut}(G)$ is quasi-oligomorphic, and $\mathrm{Out}(G)$ is profinite. $\mathrm{Out}(\mathrm{Aut}(M))$ has been studied for particular $\aleph_{0}$-categorical structures $M$ ([19, 20]) but to our knowledge this is the first time that results on $\mathrm{Aut}(G)$ have been proved at this level of generality. We believe that they will remain of interest even in case that Theorem 1.5 gets superseded by an affirmative answer to Question 1.1.
Theorem 1.8.
Let $G$ be an oligomorphic group.
  1. (A)
    If $G$ has no algebraicity, then $\mathrm{Aut}(G)$ is quasi-oligomorphic. Also $\mathrm{Out}(G)\cong N_{G}/G$ which is profinite, where $N_{G}$ is the normaliser in $\mathrm{Sym}(\omega)$.
  2. (B)
    If $G$ has finitely many essential subgroups up to conjugacy, then $\mathrm{Aut}(G)$ is quasi-oligomorphic and $\mathrm{Out}(G)$ is profinite.
Theorem 1.8 will again be proved separately for each class: (A) follows Theorem 4.7(2,3), and (B) follows from Theorem 5.8(2,3). (A) has a converse: by [8, Lemma 3.1] for every (countably based) profinite group $K$ there is an $\aleph_{0}$-categorical structure $M$ with no algebraicity such that $K\cong N_{G}/G$ where $G=\mathrm{Aut}(M)$.
It is well-known ([12, Ch. 4], also see 2.1) that there is a correspondence between oligomorphic subgroups $G$ of $\mathrm{Sym}(\omega)$ and $\aleph_{0}$-categorical structures. This allows a transfer of terminology between the two settings (already used in Def. 4.1 above). We will show that in (A), if in addition $G=\mathrm{Aut}(M)$ for a structure $M$ in a finite signature, then $\mathrm{Out}(G)$ is finite. This also has a converse: the second author and Shelah [19] have shown that for every finite group $K$ there exists a countable $\aleph_{0}$-categorical structure $M$ in a finite language such that $K\cong\mathrm{Out}(\mathrm{Aut}(M))$. In addition, $M$ has the the small index property, weak elimination of imaginaries and no algebraicity.
We leave open the following questions already discussed above.
Question 1.9.
If an oligomorphic group is $\mathfrak{G}$-finite, does it have only finitely many conjugacy classes of essential subgroups?
Question 1.10.
Is $\mathrm{Out}(G)$ always profinite for an oligomorphic group $G$?
Conventions 1.11.
By $G$ we will always denote an oligomorphic group. $\mathrm{Aut}(G)$ will denote the group of continuous automorphism of $G$. We use the group theoretic notation $g^{h}=hgh^{-1}$ for the conjugate of $g$ by $h$. By $M$ we always denote a (countably infinite) $\aleph_{0}$-categorical structure. By “definable” we will mean “00-definable” unless otherwise noted.
2. The model-theoretic framework
The language we use is mainly taken from the areas of topological groups and permutation groups. The following well-known fact can be used as a mechanism to translate from group theoretic language to model theoretic language and back.
Fact 2.1 (see [12], Theorem 4.1.4).
Let $G$ be oligomorphic. We have $G=\mathrm{Aut}(M_{G})$ for the $\aleph_{0}$-categorical relational structure $M_{G}$ that has a $n$-ary relation $P^{n}_{i}$ for each $n$-orbit of $G$. In this context one says that $M=M_{G}$ is a canonical structure for $G$; it is unique up to “renaming” the $k$-orbits, for each $k$.
The operation $G\mapsto M_{G}$ is easily seen to be Borel [18, Section 1.3]. The translation can turn the simple into the complicated. For instance, by a result of Coquand elaborated on in [1], for oligomorphic groups $G,H$, we have that $G\cong H$ iff $M_{G}$ and $M_{H}$ are bi-interpretable (an important model-theoretic notion that is somewhat tedious to define formally, see [12, Chapter 5]). However, the model theoretical language is of great importance in proving results in the setting of oligomorphic groups. For instance, it is not hard to see that bi-interpretability of $\aleph_{0}$-categorical structures is Borel [18, Section 1.3]; so topological isomorphism of oligomorphic groups is also Borel (where naively it looks merely analytical). Furthermore, using that the operation $G\mapsto M_{G}$ is Borel, one can verify that classes of groups such as the ones having no algebraicity (Def. 4.1) are Borel by looking at the corresponding classes of structures.
This section uses model theoretic language to provide results that will be important later on. For instance, we get a grip on open cosets of $G$ by representing such as coset as a pairs of imaginaries up to a certain form of interdefinability.
2.1. Imaginaries and $M^{\mathrm{eq}}$
The structure $M^{\mathrm{eq}}$ of imaginaries over $M$ goes back to Shelah [23]. It has sorts $S=D/E$ for every $k\geqslant 1$, every nonempty definable set $D\subseteq M^{k}$, and every definable equivalence relation $E$ on $D$; its elements are written as $\bar{a}/E$ for $\bar{a}\in D$. The atomic relations of $M^{\mathrm{eq}}$ are the atomic relations of the given structure $M$ (viewed as relations on the “home sort” $M$), as well as the graphs of the projections $D\to S=D/E$ for each sort $S$. For sorts $S_{i}=D_{i}/E_{i}$, $i=1,\ldots,n$, we have the product sort $\prod S_{i}=\prod D_{i}/\prod E_{i}$, where the equivalence relation $\prod E_{i}$ is defined component-wise. For detail see [24, Section X], and also the exposition of $M^{\mathrm{eq}}$ in [21, Section 16.4].
In a many-sorted structure, each definable relation is a subset of $\prod_{i<n}S_{i}$ for sorts $S_{0},\ldots,S_{n-1}$. The definable relations are given as usual via formulas in the first-order logic in the corresponding signature with variables for each sort. Extending the usual definition of algebraic/definable closure to a many-sorted structure is no problem. We note that when $M$ is $\aleph_{0}$-categorical, $\text{acl}(A)\cap S$ is finite for each finite $A\subseteq M^{\mathrm{eq}}$ and each sort $S$.
2.2. Orbital structures and profinite groups
Given an $\aleph_{0}$-categorical structure $M$, one defines a reduct $\mathcal{E}_{M}$ of $M$ such that $\mathrm{Aut}(\mathcal{E}_{M})$ is the normaliser of $\mathrm{Aut}(M)$ in $\mathrm{Sym}(M)$, the group of permutations of $M$.
Definition 2.2 (Orbital structure).
Consider the language $\{\sim_{n}:0<n<\omega\}$ where the relation symbol $\sim_{n}$ is of arity $2n$. Given an $\aleph_{0}$-categorical structure $M$, by $\mathcal{E}_{M}$ one denotes the structure in this language such that $\bar{a}\sim_{n}\bar{b}$ iff there is $g\in\mathrm{Aut}(M)$ such that $g(\bar{a})=\bar{b}$. One says that $\mathcal{E}_{M}$ is the orbital structure of $M$.
Fact 2.3.
(i) $N_{G}=\mathrm{Aut}(\mathcal{E}_{M})$. (ii) $\mathcal{E}_{M}$ is the smallest reduct $R$ of $M$ such that $\mathrm{Aut}(M)$ is a normal subgroup of $\mathrm{Aut}(R)$.
Proof.
(i) For the inclusion of left in right, let $h\in N_{G}$. Suppose that $g\cdot\bar{a}=\bar{b}$ where $g\in G$ and $\bar{a},\bar{b}\in M^{n}$. Let $g^{\prime}=hgh^{-1}\in G$. Then $g^{\prime}h\cdot\bar{a}=h\cdot\bar{b}$. So $h$ preserves the relations $\sim_{n}$. For the converse inclusion, we may assume that $M=M_{G}$. If $h\in\mathrm{Aut}(\mathcal{E}_{M})$ and $g$ preserves each relation $P^{n}_{i}$, then so does $g^{h}$. Thus $h\in N_{G}$.
(ii) Let $R$ be a reduct of $M$. Then, by (i), $\mathrm{Aut}(M)$ is normal in $\mathrm{Aut}(R)$ iff $\mathrm{Aut}(R)$ is contained in the normaliser $\mathrm{Aut}(\mathcal{E}_{M})$ of $\mathrm{Aut}(M)$, iff $\mathcal{E}_{M}$ is a reduct of $R$. ∎
Theorem 2.1.
Let $M$ be $\aleph_{0}$-categorical. Then $R=\mathrm{Aut}(\mathcal{E}_{M})/\mathrm{Aut}(M)$ is a profinite group. If $M$ is interdefinable with a structure in a finite signature then $R$ is finite.
Proof.
$M$ is interdefinable with $M_{\mathrm{Aut}(M)}$, so we may assume that for every $0<n<\omega$, $M$ has predicates $P^{n}_{1},...,P^{n}_{k(n)}$ for each of its $n$-orbits, and no other relations or functions. For $n\leqslant\omega$ let $M_{n}$ be the structure obtained after removing the predicates of arity greater than $n$, and let $\mathcal{E}_{n}:=\mathcal{E}_{M_{n}}$ be the corresponding orbital structure as in 2.2.
For $k\leqslant m\leqslant\omega$ the identity map $\mathcal{E}_{m}\to\mathcal{E}_{k}$ induces a map $q_{m,k}\colon\mathrm{Aut}(\mathcal{E}_{m})/\mathrm{Aut}(M_{m})\to\mathrm{Aut}(\mathcal{E}_{k})/\mathrm{Aut}(M_{k})$. Writing $p_{n}=q_{n+1,n}$ this yield an inverse system
$(\mathrm{Aut}(\mathcal{E}_{n})/\mathrm{Aut}(M_{n}),p_{n})_{n\in\mathbb{N}}.$
The group $\mathrm{Aut}(\mathcal{E}_{n})/\mathrm{Aut}(M_{n})$ is canonically isomorphic to a subgroup of $S_{k(n)}$ (the group of permutations of $\{1,\ldots,k(n)\}$): an isomorphism is induced by mapping $\pi\in\mathrm{Aut}(\mathcal{E}_{n})$ to the permutation $\alpha\in S_{k(n)}$ such that $\pi(P^{n}_{i})=P^{n}_{\alpha(i)}$ for each $i\leqslant k(n)$. Let
[…]
be the inverse limit in the category of topological groups, where the finite groups $\mathrm{Aut}(\mathcal{E}_{n})/\mathrm{Aut}(M_{n})$ carry the discrete topology. Every $f\in L$ can be concretely seen as a certain sequence of permutations $(\alpha_{n})_{0<n<\omega}$ where $\alpha_{n}\in S_{k(n)}$. We claim that $R\cong L$ via the continuous homomorphism $F\colon R\to L$ induced by the maps $q_{\omega,n}:R\to\mathrm{Aut}(\mathcal{E}_{n})/\mathrm{Aut}(M_{n})$. To verify this, we invoke the universal property of the inverse limit. To see that $F$ is 1-1, suppose $\pi\in\mathrm{Aut}(\mathcal{E}_{\omega})\setminus\mathrm{Aut}(M)$. Then there is $n$ and $i\leqslant k(n)$ such that $\pi(P^{n}_{i})=P^{n}_{j}$ with $i\neq j$. So the image of $\pi\mathrm{Aut}(M)$ in $\mathrm{Aut}(\mathcal{E}_{n})/\mathrm{Aut}(M_{n})$ is not equal to the identity. Thus $F(\pi\mathrm{Aut}(M))$ is not equal to the identity.
To see that $F$ is onto, given $f$ let $(\alpha_{n})_{0<n<\omega}$ be as above, and let $T_{f}$ be the theory in the language of $M$ with an additional function symbol $\pi$ with the following axioms:
  1. (1)
    $\mathrm{Th}(M)$;
  2. (2)
    $\pi$ is a permutation;
  3. (3)
    $\pi(P^{n}_{i})=P^{n}_{\alpha_{n}(i)}$, for each $n$ and each $i\leqslant k(n)$.
Note that $T_{f}$ is consistent by the Compactness Theorem, using the assumption that each $\alpha_{n}$ is induced by an automorphism $\pi_{n}$ of $\mathcal{E}_{n}$. Let $N$ be any countable model of $T_{f}$, and let $N^{\prime}$ be the reduct of $N$ to the language of $M$. Since $M$ is $\aleph_{0}$-categorical, there is an isomorphism $h\colon M\to N^{\prime}$. We have $g=h^{-1}\circ\pi^{N}\circ h\in\mathrm{Aut}(\mathcal{E}_{M})$ and $F(g\mathrm{Aut}(M))=f$. As $f\in L$ was arbitrary this shows that $F$ is onto.
The group isomorphism $F^{-1}\colon L\to R$ is clearly Borel, so it is continuous as a Borel (and hence Baire measurable) homomorphism between Polish groups; see, e.g., [11, Th. 2.3.3].
If $M$ has finite language then the inverse system above is eventually constant, and hence its inverse limit is finite. ∎
Corollary 2.2.
Let $M$ be $\aleph_{0}$-categorical. Let $R$ be a reduct of $M$ such that $\mathrm{Aut}(M)$ is a normal subgroup of $\mathrm{Aut}(R)$. Then $\mathrm{Aut}(R)/\mathrm{Aut}(M)$ is a profinite group.
Proof.
$\mathrm{Aut}(R)$ is contained $\mathrm{Aut}(\mathcal{E}_{M})$ by Fact 2.3. Since $\mathrm{Aut}(R)$ is also closed in $\mathrm{Sym}(\omega)$, we conclude that $\mathrm{Aut}(R)/\mathrm{Aut}(M)$ is profinite as a closed subgroup of $\mathrm{Aut}(\mathcal{E}_{M})/\mathrm{Aut}(M)$. ∎
2.3. Open cosets and pairs of imaginaries
[18, Section 1.2] develops a duality between oligomorphic groups $G$ and countable structures $\mathcal{M}(G)$. The domain of $\mathcal{M}(G)$ consists of the cosets of open subgroups of $G$ (note that the collection of open left cosets coincides with the open right cosets because $aU=(aUa^{-1})a$). They can be viewed as approximations to elements, because for each $g\in G$, the open cosets $A$ such that $g\in A$ form a neighbourhood basis of $g$, which is why $\mathcal{M}(G)$ is called the coarse group for $G$. The language for $\mathcal{M}(G)$ consists of a ternary relation symbol denoting that $AB\subseteq C$, the “coarse” version of the group multiplication.
Recall that $G=\mathrm{Aut}(M)$ where $M=M_{G}$ is the canonical structure for $G$ (not to be confused with the coarse group $\mathcal{M}(G)$). Note that $G$ acts naturally on $M^{\mathrm{eq}}$. Here we describe open cosets and the inclusion relation between them model-theoretically via $M^{\mathrm{eq}}$. We show in 2.11 that any two open subgroups are linked by at most finitely many cosets. This will be used to show that $\mathrm{Inn}(G)$ is closed in $\mathrm{Aut}(G)$ (Theorem 3.3). We begin with open subgroups. The following technical notion will be needed in Section 4.
Definition 2.4 ([22]).
Let $M$ be a set, $D\subseteq M^{n}$ for $n\geq 1$, and let $E$ an equivalence relation on $D$. The relation $E$ is called degenerate if for some set $\sigma\subsetneq n$, the equivalence relation $E_{\sigma,D}=\{\langle{a},{b}\rangle\in D^{2}\colon\forall i\in\sigma\,[a_{i}=b_{i}]\}$ is contained in $E$. Otherwise $E$ is called non-degenerate.
For $\alpha\in M^{\mathrm{eq}}$, let $G_{\alpha}=\{g\in G\colon g(\alpha)=\alpha\}$. The open subgroups coincide with the groups of the form $G_{\alpha}$ by the following well-known fact that is implicit in the proof of [1, Theorem 1.2]. The present form is close to Theorem 2.7 in [6].
Fact 2.5.
Let $H$ be an open subgroup of $G$. Suppose that $G_{\bar{a}}\leqslant H$ where $\bar{a}\in M^{n}$ (note that such an $\bar{a}$ exists as $H$ is open).
  1. (i)
    There are a definable relations $D\subseteq M^{n}$ and $E\subseteq M^{2n}$ such that $E$ induces an equivalence relation on $D$, and $H=G_{\bar{a}/E}$ for some $\bar{a}\in D$.
  2. (ii)
    For every $\bar{a}\in M^{n}$, only finitely many open subgroups of $G$ contain $G_{\bar{a}}$.
  3. (iii)
    If $\bar{a}$ is chosen of minimal length such that $G_{\bar{a}}\leqslant H$, then $E$ seen as an equivalence relation on $D$ can be chosen non-degenerate.
Proof.
(i) Let $D=G\cdot\bar{a}$, and let the equivalence relation $E$ on $D$ be given by
$h\bar{a}Eg\bar{a}\leftrightarrow g^{-1}h\in H.$
Since $D$ and $E$ viewed as a $2n$-ary relation are invariant under the action of $G$, they are definable. Let $\alpha=\bar{a}/E$. Then we have $G_{\alpha}=H$, because $g(\alpha)=\alpha$ iff $g(\bar{a})E\bar{a}$ iff $g\in H$.
(ii) is immediate because $M$ has only finitely many $2n$-orbits.
(iii) Assume for contradiction that $E$ is degenerate via $\sigma$. We may assume that $\sigma=\{0,\ldots,n-2\}$. Let $\bar{a}^{\prime}$ be $\bar{a}$ with the last entry removed. We claim that $G_{\bar{a}^{\prime}}\leqslant H$. For let $h\in G_{\bar{a}^{\prime}}$, and let $\bar{b}$ be the tuple $\bar{a}^{\prime}$ with $h(a_{n-1})$ appended, so that $\bar{b}=h(\bar{a})$. Then $\bar{b}\,E_{\sigma,D}\,\bar{a}$ and hence $\bar{b}\,E\,\bar{a}$ hold, so $h\in H$ by the definition of $E$. ∎
Using $M^{\mathrm{eq}}$ we can describe the containment relation between open subgroups.
Fact 2.6.
For $\alpha,\beta\in M^{\mathrm{eq}}$, we have $G_{\alpha}\subseteq G_{\beta}$Leftrightarrow $\beta\in\text{dcl}(\alpha)$.
Proof.
The implication $\Leftarrow$” is immediate.
For the implication $\Rightarrow$”, suppose $\alpha=\bar{a}/E$ and $\beta=\widetilde{b}/F$ as above. Then
$h\bar{a}Eg\bar{a}\leftrightarrow g^{-1}h\in G_{\alpha}$ and $h\widetilde{b}Fg\widetilde{b}\leftrightarrow g^{-1}h\in G_{\beta}$.
Let $R=G\cdot\langle\bar{a},\widetilde{b}\rangle$, and note that $R$ is a left invariant, and hence definable, relation. Since $G_{\alpha}\subseteq G_{\beta}$, $\beta$ is the unique equivalence class of $F$ containing $\widetilde{b}^{\prime}$ for some $\langle\bar{a}^{\prime},\widetilde{b}^{\prime}\rangle\in R$ with $\bar{a}^{\prime}\in\alpha$. Hence $\beta\in\text{dcl}(\alpha)$. ∎
Next we represent open cosets by certain pairs of imaginaries.
Notation 2.7.
For imaginaries $\alpha_{0},\alpha_{1}$ of the same sort of $M^{\mathrm{eq}}$, let
$[\alpha_{0},\alpha_{1}]=\{g\in G\colon g(\alpha_{1})=\alpha_{0}\}.$
If $[\alpha_{0},\alpha_{1}]\neq\emptyset$, then clearly $[\alpha_{0},\alpha_{1}]$ is a right coset of $G_{\alpha_{0}}$, and a left coset of $G_{\alpha_{1}}$. Also $[\alpha_{1},\alpha_{0}]=[\alpha_{0},\alpha_{1}]^{-1}$. We note that for each sort $S$, the relation $\{\langle\alpha_{0},\alpha_{1}\rangle\in S^{2}\colon[\alpha_{0},\alpha_{1}]\neq\emptyset\}$ is definable: it says that $\alpha_{0}$ and $\alpha_{1}$ are in the same orbit under the action of $G$ on $M^{\mathrm{eq}}$; now use that each orbit on $M^{\mathrm{eq}}$ is definable.
Fact 2.8.
Each open coset $A$ is of the form $[\alpha_{0},\alpha_{1}]$. Here $\alpha_{1}$ can be chosen to be any imaginary such that $A$ is a left coset of $G_{\alpha_{1}}$ (which exists by Fact 2.5(i)).
Proof.
Note that $A=hG_{\alpha_{1}}$ for any $h\in A$; now let $\alpha_{0}=h(\alpha_{1})$. ∎
The following extends Fact 2.6, given that $[\alpha,\alpha]=G_{\alpha}$.
Fact 2.9.
Suppose that $A=[\alpha_{0},\alpha_{1}]$ and $B=[\beta_{0},\beta_{1}]$ are defined. Then $A\subseteq B$ if and only if $A=\emptyset$, or $\beta_{i}\in\text{dcl}(\alpha_{i})$ for $i=0,1$ via the same formula.
Proof.
(i) The implication $\Leftarrow$” is again immediate. For $\Rightarrow$”, suppose $\alpha_{i}=\bar{a}_{i}/E$ and $\beta_{i}=\widetilde{b}_{i}/F$ are as above for $i=0,1$. Since $G_{\alpha_{1}}=A^{-1}A\subseteq B^{-1}B=G_{\beta_{1}}$, we have $\beta_{1}\in\text{dcl}(\alpha_{1})$, via a formula $\phi(\bar{x},\widetilde{y})$. We may assume $A\neq\emptyset$. Let $g\in A$, so that $g(\alpha_{1})=\alpha_{0}$. Then $g(\beta_{1})=\beta_{0}$. Hence $\beta_{0}\in\text{dcl}(\alpha_{0})$ via $\phi$. ∎
Notation 2.10.
For any sort $S$ of $M^{\mathrm{eq}}$ and $\beta\in S$, we let
$\mathrm{bidef}_{S}(\beta)=\{\alpha\in S\colon\alpha\in\mathrm{dcl}(\beta)\land\beta\in\mathrm{dcl}(\alpha)\}.$
Fact 2.11.
Let $V$ be an open subgroup of $G$. Let $S$ be a sort that is a single orbit in $M^{\mathrm{eq}}$, such that $V=G_{\beta}$ for some $\beta\in S$. Let $k=|\mathrm{bidef}_{S}(\beta)|$. Then for each open subgroup $U$, the number of right cosets of $U$ that are also left cosets of $V$ equals either 00 or $k$. In particular, setting $U=V$, we have that $|N_{G}(V):V|=k$.
Proof.
If $U$ is not conjugate to $V$ then the number of such cosets is 00. Otherwise, by 2.8, the right cosets $A$ of $U$ that are left cosets of $V$ are precisely the $[\alpha,\beta]$ where $G_{\alpha}=U$. If $A^{\prime}=[\alpha^{\prime},\beta]$ is another such coset, then $A^{\prime}=A$ iff $\alpha^{\prime}=\alpha$. So the number of cosets equals $|\mathrm{bidef}_{S}(\alpha)|$ for any such $\alpha$, which equals $k$ as $\alpha$ and $\beta$ are in the same $G$-orbit. ∎
Example 2.12.
Let $k\geqslant 2$. Let $M$ be the countably infinite structure with one permutation $\pi$, where each cycle has length $k$. Then $G\cong C_{k}\wr\mathrm{Sym}(\omega)$ where $C_{k}$ is cyclic of size $k$ and the wreath product is taken via the action of $\mathrm{Sym}(\omega)$ on $\omega$; this acts 1-transitively on $M$. Let $V=G_{b}$ where $b\in M$. Each subgroup $U$ of the form $G_{a}$ has $k$ right cosets $A_{i}$ $(i<k)$ that are left cosets of $V$: these are the $A_{i}=\{f\in G\colon\pi^{i}(a)=f(b)\}$ where $0\leqslant i<k$. Letting $a=b$ we see that the group $N_{G}(V)/V$ is cyclic of size $k$. In general, if $G$ is 1-transitive and $V=G_{b}$ then $N_{G}(V)/V$ is isomorphic to the centraliser of $G$ in $\mathrm{Sym}(\omega)$ via the map $\pi\to[\pi(b),b]$. This group can be non-abelian as a variant of the above example shows: replace the cyclic group above by a subgroup of $S_{n}$ with non-abelian centraliser.
3. $\mathrm{Aut}(G)$ as a non-Archimedean group
This section presents results of independent interest on $\mathrm{Aut}(G)$ and $\mathrm{Out}(G)$ for a general oligomorphic group $G$. A countably based topological group is non-Archimedean if it has a neighbourhood basis of $1$ consisting of open subgroups. Equivalently, it is topologically isomorphic to a closed subgroup os $\mathrm{Sym}(\omega)$. Recall the topology on $\mathrm{Aut}(G)$ from Definition 1.7. The Borel duality between $G$ and $\mathcal{M}(G)$ from [18], discussed at the beginning of Subsection 2.3, implies that there is an algebraic isomorphism $\mathrm{Aut}(G)\to\mathrm{Aut}(\mathcal{M}(G))$ that maps $\beta\in\mathrm{Aut}(G)$ to its action $A\mapsto\beta(A)$ on $\mathcal{M}(G)$. The topology given in 1.7 is obtained via this isomorphism from the topology on the non-Archimedean group $\mathrm{Aut}(\mathcal{M}(G)$ as a closed subgroup of $\mathrm{Sym}(\mathcal{M}(G))$. Thus:
Fact 3.1.
$\mathrm{Aut}(G)$ with the topology given by 1.7 is non-Archimedean.
The meet groupoid $\mathcal{W}(G)$ is a countable structure that is a variant of the coarse group $\mathcal{M}(G)$. It was briefly introduced in [18, Rem. 2.16], where it was noted that $\mathcal{M}(G)$ and $\mathcal{W}(G)$ are interdefinable via a fixed set of formulas. It is elaborated on in [17], albeit in the setting of totally disconnected, locally compact groups (where its domain consists of the compact open cosets). It will be convenient to use $\mathcal{W}(G)$, rather than $\mathcal{M}(G)$ in what follows.
Definition 3.2.
Given an oligomorphic group $G$, the domain of the meet groupoid $\mathcal{W}(G)$ consists of the open cosets, together with $\emptyset$. The groupoid product $A\cdot B$ of open cosets $A\cdot B$ is the usual product $AB$ in the group; it is defined only if $A$ is a left coset of a subgroup that $B$ is a right coset of, so that the result is again a coset. The meet is the usual intersection as a binary operation.
Note that a coset $A$ in $G$ is a right coset of the subgroup $U=A\cdot A^{-1}$ and a left coset of $V=A^{-1}\cdot A$. Given that groupoids can be seen as categories, we say that $U$ is the domain of $A$ and that $V$ is the codomain of $A$. Thus, $A\cdot B$ is defined iff the codomain of $A$ coincides with the domain of $B$.
$\mathrm{Inn}(G)$ denotes the group of automorphisms of $G$ that are given by conjugation.
Theorem 3.3.
Let $G$ be an oligomorphic group, and $N_{G}$ its normaliser in $\mathrm{Sym}(\omega)$.
  1. (i)
    $\mathrm{Inn}(G)$ is closed in $\mathrm{Aut}(G)$.
  2. (ii)
    Let $R\colon N_{G}\to\mathrm{Aut}(G)$ be the homomorphism defined by $R(\pi)(g)=\pi\circ g\circ\pi^{-1}$.
    The subgroup $\mathrm{range}(R)$ is open in $\mathrm{Aut}(G)$.
  3. (iii)
    $\mathrm{Out}(G)$ is totally disconnected and locally compact, having the profinite group $\mathrm{range}(R)/\mathrm{Inn}(G)$ as an open subgroup.
Remark 3.4.
Similar results are also obtained in work in progress of the first author with Philipp Schlicht, using the model theoretic setting of bi-interpretations. $\mathrm{Out}(G)$ corresponds to the group of self bi-interpretations of $M$ up to an appropriate congruence relation, and $\mathrm{range}(R)/\mathrm{Inn}(G)$ corresponds to the one-dimensional such bi-interpretations. So, if $M_{G}$ is interdefinable with a structure in a finite signature then $\mathrm{Out}(G)$ is countable. We expect that this work with of the first author with Schlicht will allow to describe the structure of the groups $N_{G}/G$ and $\mathrm{Out}(G)$ in terms of $\mathrm{Th}(M_{G})$ alone.
Remark 3.5.
The topology on $\mathrm{Inn}(G)$ inherited from $\mathrm{Aut}(G)$ is the expected one. For, the centre $Z=Z(G)$ of any oligomorphic group $G$ is finite because each central automorphism of $M_{G}$ is definable as a relation. Thus $Z$ is closed. We have a Polish topology on $G/Z$ by declaring $UZ/Z$ open iff $UZ$ is open in $G$. The canonical isomorphism $L\colon G/Z\to\mathrm{Inn}(G)$ is continuous. Since $\mathrm{Inn}(G)$ is Polish as a closed subgroup of $\mathrm{Aut}(G)$, $L$ is a homeomorphism by a standard result in the theory of Polish groups (see e.g., [11, 2.3.4]).
Proof of Theorem 3.3.
(i). By a fact in the theory of Polish groups (see [11, Prop. 2.2.1]), it suffices to show that $\mathrm{Inn}(G)$ is $G_{\delta}$ in $\mathrm{Aut}(G)$. Let $(A_{n})_{n\in\omega}$ be a listing of the open cosets of $G$ without repetition. Given $\Phi\in\mathrm{Aut}(G)$, we will define a set $T_{\Phi}$ of strings over some infinite alphabet which is closed under prefixes, and thus can be seen as a rooted tree. The alphabet consists of pairs of open cosets of $G$. Informally, the $n$-th level consists of all the pairs of approximations to some $g$ and $g^{-1}$ such that $gA_{i}g^{-1}=\Phi(A_{i})$ for $i<n$. Formally, we let:
$T_{\Phi}=\{\langle(B_{i},C_{i})\rangle_{i<n}\colon\Phi(A_{i})=B_{i}\cdot A_{i}\cdot C_{i}^{-1}\land\bigcap_{i<n}(B_{i}\cap C_{i})\neq\emptyset\}.$
Note that the domains of such open cosets $B_{i}$ and $C_{i}$ are determined by the condition that $\Phi(A_{i})=B_{i}\cdot A_{i}\cdot C_{i}^{-1}$. Importantly, $T_{\Phi}$ is finitely branching, because by 2.11 there are only finitely many possibilities for cosets with given domains.
Claim. $\Phi\in\mathrm{Aut}(G)$ is inner iff $T_{\Phi}$ has an infinite path.
Assuming the claim, we conclude the argument as follows. By König’s Lemma, $T_{\Phi}$ has an infinite path iff each of its levels (i.e., strings of a certain length $n$) is nonempty. Whether the $n$-th level of $T_{\Phi}$ is nonempty only depends on the values $\Phi(A_{0}),\ldots,\Phi(A_{n-1})$, so the set of such $\Phi$ is open in $\mathrm{Aut}(G)$. Thus the condition that each level is nonempty is $G_{\delta}$. So the claim shows that the subgroup $\mathrm{Inn}(G)$ is $G_{\delta}$ in $\mathrm{Aut}(G)$.
It remains to verify the claim. For the implication from left to right, suppose that there is $g\in G$ such that $\Phi(A)=gAg^{-1}$ for each open coset $A$. Suppose $A_{i}$ is a right coset of a subgroup $U_{i}$ (say $A_{i}=U_{i}h_{i}$), and a left coset of $V_{i}$ (say $A_{i}=k_{i}V_{i}$). Now let $B_{i}=gU_{i}$ and $C_{i}=gV_{i}$. Then $(\langle B_{i},C_{i})\rangle_{i\in\omega}$ is an infinite path on $T_{\Phi}$: we have $g\in\bigcap_{i<n}(B_{i}\cap C_{i})$ for each $n$, and
$\begin{array}[]{rcl}gU_{i}A_{i}V_{i}g^{-1}&=&gU_{i}h_{i}V_{i}g^{-1}\\ &=&gA_{i}V_{i}g^{-1}\\ &=&gk_{i}V_{i}g^{-1}\\ &=&gA_{i}g^{-1}.\end{array}$
For the implication from right to left, let $f$ be an infinite path on $T_{\Phi}$, and write $\langle B_{i},C_{i}\rangle=f(i)$. For each $n$, choose $g_{n}\in\bigcap_{i<n}(B_{i}\cap C_{i}^{-1})$. Given $m$, let $r=r(m)$ be the number such that
$A_{r(m)}=G_{(0,\ldots,m-1)}.$
Note that $B_{r}$ is a left coset of $G_{(0,\ldots,m-1)}$ since $B_{r}\cdot A_{r}$ is defined. If $n\geqslant r$ then for each $i<m$, we have $g_{n}(i)=g_{r}(i)$, because $g_{n}^{-1}g_{r}\in B_{r}^{-1}\cdot B_{r}$, and $B_{r}^{-1}\cdot B_{r}=G_{(0,\ldots,m-1)}$. Similarly, $g^{-1}_{n}(i)=g^{-1}_{r}(i)$ because $g_{r}g_{n}^{-1}$ in $C_{r}^{-1}\cdot C_{r}=G_{(0,\ldots,m-1)}$.
Thus $g(i):=\lim_{n}g_{n}(i)$ and $h(i):=\lim_{n}g^{-1}_{n}(i)$ exist for each $i$. Clearly the functions $g\colon\omega\to\omega$ and $h\colon\omega\to\omega$ are inverses of each other. Thus $g\in G$, being a permutation that is a pointwise limit of elements of $G$.
Given $\ell\in\omega$, we verify that $g\in B_{\ell}$. There is $m$ with $r(m)>\ell$ such that $G_{(0,\ldots,m-1)}$ is contained in the right domain $U$ of $B_{\ell}$. We have $g(i)=g_{r(m)}(i)$ for each $i<m$. Thus $g^{-1}g_{r(m)}\in G_{(0,\ldots,m-1)}\subseteq U$. Since $g_{r(m)}\in B_{\ell}$, this implies that $B_{\ell}=g_{r(m)}U=gU$, so $g\in B_{\ell}$ as required. Similarly, there is $m$ with $r(m)>\ell$ such that $G_{(0,\ldots,m-1)}$ is contained in the left domain of $C_{\ell}^{-1}$, which equals the right domain of $C_{\ell}$. Thus, arguing as before, $g\in C_{\ell}$ for each $\ell\in\omega$. Thus, $gA_{\ell}g^{-1}=B_{\ell}\cdot A_{\ell}\cdot C^{-1}_{\ell}=\Phi(A_{\ell})$ for each $\ell$, so $\Phi$ is inner.
(ii). We will use Notation 2.7, where $M$ is the canonical structure for $G$ as in 2.1 (with domain $\omega$). Note that for $a,b,c\in\omega$, if $[b,a]\supseteq[c,a]\neq\emptyset$ then $b=c$: By 2.9, there is a definable function $\eta$ on $\omega$ such that $b=\eta(c)$ and $a=\eta(a)$. Such a function commutes which each $g\in G$. Since $c=g(a)$ for some $g\in G$, we have $b=\eta(g(a))=g(\eta(a))=c$.
Let $a_{1},\ldots,a_{n}$ represent the $1$-orbits of $G$. It suffices to show that the open subgroup $\{\Phi\in\mathrm{Aut}(G)\colon\,\Phi(G_{a_{i}})=G_{a_{i}}\text{ for each }i\}$ is contained in $\mathrm{range}(R)$. Suppose that $\Phi$ is in this subgroup. Since $\Phi$ fixes each subgroup $G_{a_{i}}$, for each coset $D=[b,a_{i}]$, $\Phi(D)$ is also a left coset of $G_{a_{i}}$. Hence, by 2.8, $\Phi(D)$ can be written in the form $[d,a_{i}]$. As noted above, $d$ is uniquely determined by $b$. Define a function $\pi_{\Phi}\colon\omega\to\omega$ by
$\pi_{\Phi}(b)=d\text{ if }\Phi([b,a_{i}])=[d,a_{i}],\text{ where }[b,a_{i}]\neq\emptyset.$
Clearly $\pi_{\Phi}$ and $\pi_{\Phi^{-1}}$ are inverses. So $\pi_{\Phi}\in\mathrm{Sym}(\omega)$. The following establishes (ii).
Claim. $\pi_{\Phi}\in N_{G}$ and $R(\pi_{\Phi})=\Phi$.
Write $\pi=\pi_{\Phi}$. To verify the claim, we first show that $\Phi([r,s])=[\pi(r),\pi(s)]$ for each $r,s\in\omega$ such that $[r,s]\neq\emptyset$. Note that by hypothesis on $\Phi$ and since $G_{a_{i}}=[a_{i},a_{i}]$, we have $\pi(a_{i})=a_{i}$ for each $i$. Also note that $\Phi([b,c])^{-1}=\Phi([c,b])$ for each $b,c\in\omega$. Let now $i$ be such that $r$ and $s$ are in the $1$-orbit of $a_{i}$. We have $[r,s]=[r,a_{i}]\cdot[a_{i},s]=[r,a_{i}]\cdot[s,a_{i}]^{-1}$. So
$\Phi([r,s])=\Phi([r,a_{i}])\cdot\Phi([s,a_{i}])^{-1}=[\pi(r),a_{i}]\cdot[\pi(s,a_{i}]^{-1}=[\pi(r),\pi(s)].$
Next, for each $h\in G$ we have $\{h\}=\bigcap_{s\in\omega}hG_{s}$, and $hG_{s}=[h(s),s]$ using Notation 2.7. Then
$\{\Phi(g)\}=\bigcap_{s\in\omega}\Phi([g(s),s])=\bigcap_{s\in\omega}[\pi(g(s)),\pi(s)]=\bigcap_{t\in\omega}[\pi(g(\pi^{-1}(t))),t]=\{g^{\pi}\}$,
as required. This shows the claim.
(iii). The kernel of $R$ is the centraliser $C_{G}$ of $G$ in $\mathrm{Sym}(\omega)$, which consists of the permutations that are definable in $M$. Thus $C_{G}$ is a finite normal subgroup of $N_{G}$. So $GC_{G}$ is a normal subgroup of $N_{G}$. $GC_{G}$ is also closed in $\mathrm{Sym}(\omega)$ because $G$ has finite index in it. Since $C_{G}$ is the kernel of $R$ and $R(G)=\mathrm{Inn}(G)$, we have $R^{-1}(\mathrm{Inn}(G))=GC_{G}$. Thus $\mathrm{range}(R)/\mathrm{Inn}(G)$ is topologically isomorphic to $N_{G}/GC_{G}$. By 2.1 $N_{G}/G$ is compact. So $N_{G}/GC_{G}$ is also compact as a topological quotient. ∎
4. Groups with no algebraicity
A countable $\aleph_{0}$-categorical structure $M$ has no algebraicity if $\mathrm{acl}(A)=A$ for each $A\subseteq M$; equivalently, $\mathrm{Aut}(M)$ has no algebraicity in the sense of 4.1. The following will be needed to prove Theorem 4.7(2).
Fact 4.1.
If $M$ has no algebraicity, then the centraliser of $G=\mathrm{Aut}(M)$ in $\mathrm{Sym}(\omega)$ is trivial. In particular, $\mathrm{Aut}(M)$ has trivial center.
Proof.
Any permutation $h$ in the centraliser of $G$ is invariant under the action of $G$ on $M$, and hence definable as a relation. So $h$ is the identity. ∎
We review some results due to M. Rubin [22]. Fix a countable $\aleph_{0}$-categorical structure $M$.
Definition 4.2 ([22, Definition 2.9]).
Let $N$ be a substructure of $M^{\mathrm{eq}}$. We say that $N$ is normal if its universe is the union of orbits of $(\mathrm{Aut}(M),M^{\mathrm{eq}})$. We say that $N$ is a dense substructure of $M^{\mathrm{eq}}$ if it is normal and for every $b\in\mathrm{M}^{\mathrm{eq}}$ there is a finite subset $A\subseteq N$ such that $b$ is definable from $A$ in $M^{\mathrm{eq}}$. (Note that this property is called “regular” in [22].)
Notation 4.3 ([22, Definition 2.5]).
For a structure $M$ and a complete $n$-type $p$ in $M$, by $M_{p}=p(M)$ one denotes $\{\bar{a}\in M^{n}:\mathrm{tp}_{M}(\bar{a})=p\}$.
Fact 4.4 ([22, Lemma 2.10]).
Let $N$ be a normal substructure of $M^{\mathrm{eq}}$ such that $(\mathrm{Aut}(M),N)$ has finitely many orbits, is faithful, and closed. Then $N$ is a dense substructure of $M^{\mathrm{eq}}$.
Fact 4.5 ([22, Lemma 2.11]).
Suppose that $M$ has no algebraicity, $N\subseteq M^{\mathrm{eq}}$ has no algebraicity, and $N$ is normal and dense. Further suppose that if the sort $M_{p}/E$ is contained in $N$, where $p$ is a complete $n$-type and $E$ is an equivalence relation on $p(M^{n})$, then $E$ is non-degenerate in the sense of 2.4. Then $M=N$.
The following should be seen as a corollary to Rubin’s results. See [12, Chapter 5] for a formal definition of bi-interpretability.
Theorem 4.6.
Let $M$ and $N$ be countable $\aleph_{0}$-categorical structures with no algebraicity. If $M$ and $N$ are bi-interpretable then they already are bi-definable.
Proof.
We may assume that the domain of $N$ as interpreted in $M$ is of the form $D/E$, where $D\subseteq M^{k}$ is definable and $E$ is a definable equivalence relation on $D$. We claim that we can assume $E$ is non-degenerate (Def. 2.4): otherwise, say $E$ contains the equivalence relation $E_{\{0,\ldots,k-2\},D}$ expressing that two $k$-tuples over $D$ agree on the first $k-1$ elements. Let $D^{\prime}=p(D)$ where $p$ is the definable projection $M^{k}\to M^{k-1}$ that erases the last component. Then $D^{\prime}$ is definable, and $p$ yields an $M$-definable bijection $D/E\to D^{\prime}/p(E)$. So we can replace $N$ with the structure $N^{\prime}$ that has domain $D^{\prime}/p(E)$.
The result will follow via 4.5 once we can show that $N$ is a dense subset of $M^{\mathrm{eq}}$. Since $N$ is interpreted in $M$, $(\mathrm{Aut}(M),N)$ has only finitely many orbits. As $M$ and $N$ are bi-interpretable we have that $(\mathrm{Aut}(M),N)$ is faithful and closed. As $N$ is a normal substructure of $M^{\mathrm{eq}}$, it now suffices to invoke 4.4. ∎
Recall that $\mathrm{Aut}(G)$ denotes the group of continuous automorphisms of $G$. The orbital structure $\mathcal{E}_{M}$ is defined in 2.2. The following establishes 1.5(A) and 1.8(A).
Theorem 4.7.
  1. (1)
    The class of oligomorphic groups with no algebraicity is smooth.
  2. (2)
    Let $G=\mathrm{Aut}(M)$ where $M$ is an $\aleph_{0}$-categorical structure with no algebraicity. Then $\mathrm{Aut}(G)\cong\mathrm{Aut}(\mathcal{E}_{M})$.
  3. (3)
    $\mathrm{Out}(G)\cong\mathrm{Aut}(\mathcal{E}_{M})/G,$ and so $\mathrm{Out}(G)$ is a profinite group.
Proof.
(1) Recall (2.1) that $M_{G}$ denotes the canonical structure $M$ such that $\mathrm{Aut}(M)=G$, and that the operation $G\mapsto M_{G}$ is Borel [18, Secton 1.3]. It is clearly a Borel property of structures with domain $\omega$ to have no algebraicity. So the class of groups with no algebraicity is Borel. By the above mentioned result of Coquand [1], for oligomorphic groups $G,H$, we have that $G\cong H$ iff $M_{G}$ and $M_{H}$ are bi-interpretable. Since $G\mapsto M_{G}$ is Borel, for smoothness of the class it suffices to show that the relation of bi-interpretability between $\aleph_{0}$-categorical structures with no algebraicity is smooth. By 4.6 the latter is equivalent to bi-definability of such structures. Bi-definability of $\aleph_{0}$-categorical structures with domain $\omega$ is the same as conjugacy of their automorphism groups, which is smooth by [18, Prop 4.2].
(2) Fix $a_{1},...,a_{k}\in M$ as representatives of the $1$-orbits of $G$ acting on $M=M_{G}$.
Given $\beta\in\mathrm{Aut}(G)$, we will define a corresponding map $L_{\beta}\colon M\to N$ where $N\subset M^{\mathrm{eq}}$ is a submodel given by finitely many sorts. Our goal is to show that $N=M$ and hence $L_{\beta}\colon M\to M$.
The construction of $L_{\beta}$ is similar to the one in the proof of [1, Thm. 1.2]. For $i\leqslant k$ let $b_{i}\in M^{\mathrm{eq}}$ be such that $G_{b_{i}}=\beta(G_{a_{i}})$. By 2.5 we may assume that $b_{i}$ is in a sort $D_{i}/E_{i}$ where $E_{i}$ is nondegenerate. Define
$L_{\beta}(g\cdot a_{i})=\beta(g)\cdot b_{i}.$
(This is well-defined because $g\cdot a_{i}=h\cdot a_{i}$ iff $g^{-1}h\in G_{a_{i}}$ iff $\beta(g)\cdot b_{i}=\beta(h)(b_{i})$.)
The permutation groups $(G,M)$ and $(G,N)$ are isomorphic via $\langle g,a\rangle\mapsto\langle\beta(g),L_{\beta}{(A)}\rangle$. Hence, $(G,N)$ has only finitely many orbits and is faithful and closed. Thus, by 4.4 $N$ is a dense substructure of $M^{\mathrm{eq}}$. Further, $N$ has no algebraicity since this holds for $M$, and is a property that can be seen from $(G,M)$, as indicated in Definition 4.1. Using that the sorts of $N$ are non-degenerate, by 4.5, $N=M$, so that $L_{\beta}\colon M\to M$.
We now verify that we have a (topological) isomorphism $L\colon\mathrm{Aut}(G)\to\mathrm{Aut}(\mathcal{E}_{M})$ given by $\beta\mapsto L_{\beta}$. We mainly employ the two conditions (A) and (B) below:
  1. (a)
    $G_{L_{\beta}(A)}=\beta(G_{a})$ for each $a\in M$.
  2. (b)
    For each $k\geqslant 1$ and $\bar{a},\bar{b}\in M^{k}$, we have $G_{\bar{a}}=G_{\bar{b}}$ iff $\bar{a}=\bar{b}$. In particular, $h\cdot\bar{a}=\bar{c}$ iff $hG_{\bar{a}}h^{-1}=G_{\bar{c}}$ for each $h\in\mathrm{Aut}(\mathcal{E}_{M})$ (the normaliser of $G$ in $\mathrm{Sym}(\omega)$).
Condition (a) is easy to check from the definitions. Condition (b) holds because of 2.6 and since $M$ has no algebraicity. From (a) we have that $L(\beta^{-1})=(L_{\beta})^{-1}$. In particular, $L_{\beta}$ is a permutation of $M$, and $\beta\to L_{\beta}$ preserves inverses.
We verify that $L_{\beta}$ is an automorphism of $\mathcal{E}_{M}$. Fix $n$ with $0<n<\omega$. Using the notation of the definition of $\mathcal{E}_{M}$ in 2.2, we need to show that
$\bar{a}\sim_{n}\bar{b}$ iff $L_{\beta}(\bar{a})\sim_{n}L_{\beta}(\bar{b})$, for each $\bar{a},\bar{b}\in M^{n}$.
If $\bar{a}\sim_{n}\bar{b}$, there is $g\in G$ such that $g(a_{i})=b_{i}$ for all $i\in[1,n]$. Clearly $gG_{a_{i}}g^{-1}=G_{b_{i}}$. As $\beta\in\mathrm{Aut}(G)$, this implies $\beta(g)\beta(G_{a_{i}})\beta(g)^{-1}=\beta(G_{b_{i}})$ and so by (a) and (b), $\beta(g)\cdot L_{\beta}(a_{i})=L_{\beta}(b_{i})$. Thus $\beta(g)\in G$ is a witness of the fact that $L_{\beta}(\bar{a})\sim_{n}L_{\beta}(\bar{b})$, as desired. The converse implication is proved analogously, using $\beta^{-1}$ instead of $\beta$.
To see that $L$ is a group homomorphism, we note that for $\gamma,\beta\in\mathrm{Aut}(G)$, by (a) we have $G_{L_{\gamma\circ\beta}(A)}=G_{L_{\gamma}(L_{\beta}(A))}$; this suffices by (b) and since $L$ preserves inverses.
To see that $L$ is onto, given $r\in\mathrm{Aut}(\mathcal{E}_{M})$, define $\beta\in\mathrm{Aut}(G)$ by $\beta(g)=rgr^{-1}$; then $G_{L_{\beta}(A)}=\beta(G_{a})=rG_{a}r^{-1}=G_{r(A)}$ for each $a\in M$, so that $L_{\beta}=r$.
To show that $L$ is 1-1, let
$K=\mathrm{ker}(L)=\{\beta\in\mathrm{Aut}(G):\beta(G_{a})=G_{a},\text{ for all }a\in M\}.$
Note that $K\cap\mathrm{Inn}(G)=\{e\}$, because, for every $g\in G$, if conjugation by $g$ is in $K$ then for every $a\in M$ we have $G_{g(A)}=gG_{a}g^{-1}=G_{a}$, so that $g$ is the identity by (a). As $\mathrm{Inn}(G)$ is always normal in $\mathrm{Aut}(G)$, we conclude that $[K,\mathrm{Inn}(G)]=\{e\}$. An elementary group theoretic fact recalled as 4.8 below, together with 4.1, now shows that the kernel $K$ of $L$ is trivial.
It remains to show that $L$ is continuous with respect to the Polish topology on $\mathrm{Aut}(G)$ discussed in Section 3; in this case, $L$ is a homeomorphism (see, e.g., [11, 2.3.4]). A system of neighbourhoods of the identity in $\mathrm{Aut}(\mathcal{E}_{M})$ is given by the subgroups $\{r\colon\bigwedge_{i=1}^{n}\,r(a_{i})=a_{i}\}$, where $a_{1},\ldots,a_{n}\in M$. By (a) and (b), the pre-image under the map $L$ of such a group is $\{\beta\colon\,\bigwedge_{i=1}^{n}\beta(G_{a_{i}})=G_{a_{i}}\}$, which is open in $\mathrm{Aut}(G)$.
(3) The isomorphism $L:\mathrm{Aut}(G)\to\mathrm{Aut}(\mathcal{E}_{M})$ induces an isomorphism between the closed normal subgroups $\mathrm{Inn}(G)$ and $G=\mathrm{Aut}(M)$: For $r\in\mathrm{Aut}(\mathcal{E}_{M})$, the pre-image $L^{-1}(r)$ is the map $g\mapsto g^{r}$. Then $r\in\mathrm{Aut}(M)$ iff this map is an inner automorphism. The topological group $\mathrm{Aut}(\mathcal{E}_{M})/\mathrm{Aut}(M)$ is profinite by 2.1. ∎
Lemma 4.8.
Let $G$ be a group, and suppose that an automorphism $\beta$ of $G$ centralises $\mathrm{Inn}(G)$. Then for each $g\in G$ we have $\beta(g)g^{-1}\in Z(G)$.
Proof.
Let $\hat{g}\in\mathrm{Aut}(G)$ denote conjugation by $g$. For each $h\in G$ we have
$(\beta(h))^{\beta(g)}=(\beta\circ\hat{g})(h)=(\hat{g}\circ\beta)(h)=(\beta(h))^{g}.$
Thus $\beta(g)g^{-1}\in Z(G)$ as required. (The converse implication also holds.) ∎
In the remainder of the section we make some remark related to the proof of 4.7. The following fact is a well-known observation of Hrushovski; see [7, Ch. 1, Lemma 1.1.1] where it is proved in the language of permutation groups. In the subsequent remark we use the proof to show that the hypothesis that both $M$ and $N$ have no algebraicity is necessary in 4.6.
Fact 4.9.
Let $M$ be a countable $\aleph_{0}$-categorical structure. There is a countable $\aleph_{0}$-categorical structure $N$ such that $M$ and $N$ are bi-interpretable, and $N$ has only one $1$-orbit.
Proof.
Let $0<n<\omega$ be the number of $1$-orbits of $M$. We may replace $M$ by the canonical structure for $\mathrm{Aut}(M)$, by naming its $n$-orbits with predicates of arity $n$. In particular, let $P_{1},...,P_{n}$ be the predicates naming the $1$-orbits. Let
$D$ $=$ $\bigcup_{\sigma\in S_{n}}\prod_{i=1}^{n}P_{\sigma(i)}$
$E$ $=$ $\{\langle\bar{x},\bar{y}\rangle\in D^{2}\colon\exists\sigma\in S_{n}\forall i\,[x_{i}=y_{\sigma(i)}]\}.$
The structure $N$ induced by $M^{\mathrm{eq}}$ on $D/E$ is what we are looking for. Notice that $N$ is a dense subset of $M^{\mathrm{eq}}$ in the sense of 4.2. ∎
Remark 4.10.
In the context of the proof of 4.9, if $M$ has more than one $1$-orbit, then $N$ has algebraicity, i.e., there is finite $A\subseteq N$ and $a\in N\setminus A$ s.t. $a\in\mathrm{acl}_{N}(A)$.
Proof.
To save on notation, we show this for the case that $M$ has only two distinct infinite $1$-orbits, denoted $P_{1}$ and $P_{2}$. Given $\langle a,b\rangle\in D$, we denote its equivalence class by $\{a,b\}$. On $N$ we consider the undirected graph with the edge relation
$\{a,b\}R\{c,d\}$ iff |$|\{a,b\}\cap\{c,d\}|=1$.
This relation is definable because $N$ inherits all the definable relations from $M^{\mathrm{eq}}$. Fix $x\in P_{1}$ and $y\in P_{2}$. Furthermore, fix pairwise distinct $x_{i}\in P_{1}\setminus\{x\}$ and $y_{i}\in P_{2}\setminus\{y\}$, for $i=1,2$. Define $a_{i}=\{x,y_{i}\}$ and $b_{i}=\{x_{i},y\}$. Now $c=\{x,y\}$ is the unique vertex connected to each vertex in $A=\{a_{1},a_{2},b_{1},b_{2}\}$. Thus $c\in\mathrm{dcl}_{N}(A)\setminus A$. ∎
5. Groups with finitely many essential subgroups up to conjugacy
5.1. General result involving a property $P$ of open subgroups
Recall that by default, isomorphisms between topological groups are assumed to be topological.
Definition 5.1.
  1. (a)
    A property $P$ of open subgroups of closed subgroups $G$ of $\mathrm{Sym}(\omega)$ is Borel if $\{\langle G,U\rangle\colon U\text{ has }P\text{ in }G\}$ is Borel.
  2. (b)
    We say that $P$ is isomorphism invariant if any (topological) isomorphism between groups $\alpha\colon G\to H$ preserves $P$.
  3. (c)
    We call an invariant Borel property $P$ nice for $G$ if each open subgroup of $G$ contains a finite intersection of $P$-subgroups.
Condition (a) tends to hold for natural properties of open subgroups. Condition (b) rules out the property “being of the form $G_{a}$ for some $a\in\omega$” (which satisfies the other two conditions), for this depends on $G$ as a group of permutations. Condition (c) says that the $P$-subgroups form a sub-basis of neighbourhoods of the identity.
Given a property $P$ that is nice for an oligomorphic group $G$, we will define a countable structure $\mathcal{C}^{P}_{G}$, dual to $G$, that is a “trimmed-down” version of the meet groupoid $\mathcal{W}(G)$ of Definition 3.2. It will play the role of $\mathcal{E}_{M}$ in the proof of Theorem 4.7. The domain of $\mathcal{C}^{P}_{G}$ consists of the cosets of $P$-subgroups, and in place of the meet operation we merely have non-disjointness relations. A structure of form $\mathcal{C}^{P}_{G}$ is similar to the structures in [20, Def. 2.15] used in the particular case of groups of the form $\mathrm{Aut}(M)$ where $M$ has w.e.i. However, to get the duality, the partial product operation is needed; this operation was absent in [20, Def. 2.15].
Definition 5.2.
Suppose that a property $P$ of open subgroups is nice for $G$.
The domain of $\mathcal{C}^{P}_{G}$ is $\{gH:H\text{ is a $P$-subgroup of $G$}\text{ and }g\in G\}$. On this domain the following are given.
  1. (a)
    Unary functions $\mathrm{Dom}$ and $\mathrm{Cod}$ (which stand for “domain” and “codomain”) such that $\mathrm{Dom}(B)=U$ if $B$ is a right coset of the subgroup $U$, and $\mathrm{Cod}(B)=V$ if $B$ is a left coset of the subgroup $V$.
  2. (b)
    An $n$-ary relation $I_{n}$ holds of $(B_{1},\ldots,B_{n})$ if $B_{1}\cap\ldots\cap B_{n}\neq\emptyset$;
  3. (c)
    A ternary relation that holds of $(A,B,C)$ if $\mathrm{Cod}(A)=\mathrm{Dom}(B)$ and $AB=C$. As before in the case of $\mathcal{W}(G)$, we write $A\cdot B=C$ for this.
We will omit the superscript $P$ in $\mathcal{C}^{P}_{G}$ if $P$ is understood from the context. The next lemma establishes the duality between oligomorphic groups for which property $P$ is nice, and the corresponding structures.
Lemma 5.3.
Let $G,H$ be oligomorphic groups such that property $P$ is nice for $G$ and for $H$. If $\alpha$ is an isomorphism $G\to H$, let $R(\alpha)\colon\mathcal{C}_{G}\to\mathcal{C}_{H}$ be the map given by $B\mapsto\alpha(B)$.
  1. (1)
    $R(\alpha)$ is an isomorphism of the structures $\mathcal{C}_{G}$ and $\mathcal{C}_{H}$.
  2. (2)
    Each isomorphism $\theta:\mathcal{C}_{G}\to\mathcal{C}_{H}$ is of the form $R(\alpha)$ for a unique $\alpha$.
  3. (3)
    In the case $G=H$, the map $R$ witnesses that $\mathrm{Aut}(G)\cong\mathrm{Aut}(\mathcal{C}_{G})$.
Proof.
We have $G=\mathrm{Aut}(M)$ and $H=\mathrm{Aut}(N)$ where $M,N$ are the canonical structures for $G$ and $H$, respectively.
(1) Given that $\alpha$ is a topological isomorphism, it only remains to check that the relations $\mathrm{Dom}$ and $\mathrm{Cod}$ are preserved in both directions by $R(\alpha)$. For each elements $B,U,V$ in a structure of the form $\mathcal{C}_{L}$, we have $\mathrm{Dom}(B,U)$ iff $B\cdot B^{-1}=U$, and $\mathrm{Cod}(B,V)$ iff $B^{-1}\cdot B=V$. This suffices because by its definition, $R(\alpha)$ preserves inversion.
(2) Let $(U_{i})_{i<\omega}$ list the subgroups in $\mathcal{C}_{G}$, and let $V_{i}:=\theta(U_{i})$. Then $(V_{i})_{i<\omega}$ lists the subgroups in $\mathcal{C}_{H}$. We define an isomorphism $\alpha\colon G\to H$ such that $R(\alpha)=\theta$. Given $g\in G$, first we define the value $\alpha(g)$, which will be a function $\omega\to\omega$. Let $A_{i}=gU_{i}$ and $B_{i}=\theta(A_{i})$. For each $n$ we have $\mathcal{C}_{G}\models I_{n}(A_{1},\ldots,A_{n})$. So, since $\theta$ is an isomorphism, there is $r_{n}\in\bigcap_{i<n}B_{i}$. Since $\mathrm{Cod}(B_{i},V_{i})$ holds in $\mathcal{C}_{H}$, we have $r_{n}V_{i}=B_{i}$ for each $i<n$. Since $P$ is nice for $H$, for each $b\in N$ there is a finite set $I\subset\omega$ such that $\bigcap_{i\in I}V_{i}\subseteq H_{b}$. So
$\alpha(g)(B)\colon=\lim_{n}r_{n}(B)$
exists for each $b$.
We verify that $\alpha$ is a homomorphism $G\to H$. For the rest of this proof, to avoid an awkward buildup of parentheses due to iterated function applications, we use the following notation: instead of $f(x)$ we write $f.x$ (read as “$f$ of $x$”). So $f(x)(z)$ becomes $f.x.z$.
First we verify that $\alpha.e_{G}=e_{H}$. For $g=e_{G}$, in the definition of $\alpha.g$ we have $A_{i}=U_{i}$ and so $B_{i}=V_{i}$; thus $r_{n}\in\bigcap_{i<n}V_{i}$. For any $x\in\omega$, we have $\alpha.e_{G}.x=\lim_{n}r_{n}.x$. Since $r_{n}\in\bigcap_{i<n}V_{i}$, this implies $\alpha.e_{G}.x=x$.
Next we verify that for each $g,h\in G$, we have $(\alpha.h)\circ(\alpha.g)=\alpha.(hg)$. Let $U^{\prime}_{i}=gU_{i}g^{-1}$ and $V^{\prime}_{i}=\theta(U^{\prime}_{i})$. Also let $A_{i}=\theta(gU_{i})$ and $B_{i}=\theta(hU_{i}^{\prime})$. Given $x\in\omega$ write $y=\alpha.g.x$. Let $n$ be so large that $\bigcap_{i<n}V_{i}\subseteq H_{x}\cap H_{y}$. Fix $r\in\bigcap_{i<n}A_{i}$. Then $r.x=y$. Using that the subgroups with $P$ are closed under conjugation, let $m\geqslant n$ be so large that each $U_{i}$, $i<n$, is of the form $U_{j}^{\prime}$ for some $j<m$. Let $s\in\bigcap_{i<m}B_{i}$. Then $s\in\bigcap_{i<n}\theta(hU_{i})$, and hence $s.y=\alpha.h.y$.
Since $hU_{i}^{\prime}\cdot gU_{i}=hgU_{i}$ holds in $\mathcal{C}_{G}$ and $\theta:\mathcal{C}_{G}\to\mathcal{C}_{H}$ is an isomorphism, for each $i$, $B_{i}\cdot A_{i}=\theta(hgU_{i})$ holds in $\mathcal{C}_{H}$. Thus $s\circ r\in\bigcap_{i<n}\theta(hgU_{i})$. By the choice of $n$ and since $\mathrm{Cod}(\theta(hgU_{i}))=V_{i}$, this shows that $(\alpha.h)\circ(\alpha.g)(x)=s\circ r(x)=\alpha.(hg).x$.
Letting $h=g^{-1}$, we infer that $\alpha(g)\in\mathrm{Sym}(\omega)$. Then clearly $\alpha(g)$ is in $H$ because $H$ is a closed subset of $\mathrm{Sym}(\omega)$.
Exchanging the roles of $\mathcal{C}_{G}$ and $\mathcal{C}_{H}$ and using $\theta^{-1}\colon\mathcal{C}_{H}\to\mathcal{C}_{G}$, we find an inverse of $\alpha$. Thus $\alpha:G\to H$ is a group isomorphism. To see that $\alpha$ is continuous at $e_{G}$ (and hence continuous), it suffices to show that $\alpha^{-1}(H_{x})$ contains an open subgroup, for each $x\in N$. As before let $n$ be such that $\bigcap_{i<n}V_{i}\subseteq H_{x}$; then $\bigcap_{i<n}U_{i}$ is the required subgroup. Similarly, $\alpha^{-1}$ is continuous.
(3) For each $\alpha,\beta\in\mathrm{Aut}(G)$, we have $R(\alpha\beta^{-1})=R(\alpha)R(\beta)^{-1}$ by the definition of $R$, so $R$ is an isomorphism of groups. To see that $R$ is continuous (and hence a homeomorphism using [11, 2.3.4]), note that a basic neighbourhood of the identity in $\mathrm{Aut}(\mathcal{C}_{G})$ has the form $\{\theta\colon\bigwedge_{i=1}^{n}\theta(A_{i})=A_{i}\}$, where $A_{1},\ldots,A_{n}\in\mathcal{C}_{G}$. Its preimage under $R$ is open in $\mathrm{Aut}(G)$ according to Definition 1.7. ∎
By 5.3(3), $\mathrm{Aut}(G)$ acts on $\mathcal{C}_{G}$ by automorphisms, namely, $\Phi\cdot A=\Phi(A)$ where $\Phi\in\mathrm{Aut}(G)$ and $A\in\mathcal{C}_{G}$. Under an additional hypothesis, the subaction of $\mathrm{Inn}(G)$ on $\mathcal{C}_{G}$ is oligomorphic.
Lemma 5.4.
Let $G$ be an oligomorphic group such that property $P$ is nice for $G$, and there are only finitely many conjugacy classes of $P$-subgroups. Then the canonical action by automorphisms of $\mathrm{Inn}(G)$ on $\mathcal{C}_{G}$ is oligomorphic.
Proof.
By [25, Theorem 2.4] it suffices to show that the action has finitely many $1$-orbits, since $G$ as an oligomorphic group is Roelcke precompact. Note that the $\mathrm{Inn}(G)$-orbit of an element of $\mathcal{C}_{G}$ is firstly given by a pair $U,V$ of conjugate $P$-subgroups; by hypothesis they are of the form $G_{\alpha}$ and $G_{\beta}$ for $\alpha,\beta$ both elements of one of finitely many 1-transitive sorts $S$ of $M^{\mathrm{eq}}$. There are only finitely many $M^{\mathrm{eq}}$ orbits of such pairs $\alpha,\beta$. Secondly, by 2.11 the orbit of such a coset is now given by $k$ many possibilities, where $k$ only depends on $S$. ∎
Now let $\mathcal{C}^{+}_{G}$ the expansion of $\mathcal{C}_{G}$ obtained naming the $n$-orbits of the action of $\mathrm{Inn}(G)$ on $\mathcal{C}_{G}$ by predicates, for all $n\geq 1$. Note that $\mathcal{C}_{G}$ is a reduct of $\mathcal{C}^{+}_{G}$. By the above we have:
Fact 5.5.
Suppose that $G$ is as in 5.4. Then $\mathcal{C}^{+}_{G}$ is $\aleph_{0}$-categorical, and so $\mathcal{C}_{G}$ is $\aleph_{0}$-categorical as well.
Definition 5.6.
Given an invariant Borel property of open subgroups $P$, let $\mathcal{B}_{P}$ be the collection of oligomorphic groups $G$ such that $P$ is nice for $G$, and there are finitely many conjugacy classes of $P$-subgroups.
Fact 5.7.
The class $\mathcal{B}_{P}$ is Borel.
Proof.
The structures with domain $\omega$ in a fixed countable signature $L$ can be seen as a standard Borel space $\mathrm{Mod}(L)$ [11, Section 3.5]. By [15, Section 3.3], there is a Borel function taking an oligomorphic group $G$ to a bijection $\nu_{G}\colon\mathcal{\omega}\to\mathrm{dom}(\mathcal{M}(G))$; this uses a result of Lusin-Novikov in the version of [14, 18.10]. Via this bijection we can view $\mathcal{M}(G)$ as a structure with domain $\omega$ in such a way that the map $G\mapsto\mathcal{M}(G)$ is Borel. The structures $\mathcal{M}(G)$ and the meet groupoid $\mathcal{W}(G)$ are interdefinable in a uniform way [18, Remark 2.16], so the map $G\to\mathcal{W}(G)$ is also Borel in this sense. Then, since $P$ is a Borel property, the map $G\mapsto\mathcal{W}^{P}(G)$ is Borel, where $\mathcal{W}^{P}(G)$ is $\mathcal{W}(G)$ extended by a unary predicate $L_{P}$ saying that a subgroup satisfies $P$. That $P$ is nice for $G$ can be expressed by a simple $L_{\omega_{1},\omega}$-sentence about $\mathcal{W}^{P}(G)$, so this is Borel. Conjugacy of subgroups $U,V$ can be expressed in $\mathcal{W}(G)$ via the formula $\exists A[U=A\cdot A^{-1}\land V=A^{-1}\cdot A]$. So there is another $L_{\omega_{1},\omega}$-sentence saying that there are finitely many conjugacy classes of $P$-subgroups. ∎
Theorem 5.8.
Let $P$ be an invariant Borel property of open subgroups.
  1. (1)
    Topological isomorphism on the class $\mathcal{B}_{P}$ is smooth.
  2. (2)
    $\mathrm{Aut}(G)\cong\mathrm{Aut}(\mathcal{C}_{G})$ for each $G\in\mathcal{B}_{P}$, so $\mathrm{Aut}(G)$ is quasi-oligomorphic.
  3. (3)
    $\mathrm{Aut}(G)/\mathrm{Inn}(G)$ is profinite for each $G\in\mathcal{B}_{P}$.
Proof.
(1) The domain of the structure $\mathcal{C}^{P}_{G}$ can be first-order defined in the extended structure $\mathcal{W}^{P}(G)$ from the proof of 5.7 as the set of cosets of subgroups satisfying the property $L_{P}$. The relations and operation on $\mathcal{C}^{P}_{G}$ can be defined in $\mathcal{W}(G)$ in a uniform way. Viewing $\mathcal{W}(G)$ as a structure with domain $\omega$ as above, this shows that the map $G\to\mathcal{C}^{P}_{G}$ is Borel. By 5.3 and 5.5, the Borel map $G\mapsto\mathrm{Th}(\mathcal{C}^{P}_{G})$ now establishes smoothness of $\cong$ on $\mathcal{B}$.
(2) follows from 5.3 via the map $R$ defined there, namely, $R.\alpha.A=\alpha(A)$.
(3) By definition of $\mathcal{C}_{G}^{+}$, the closure of $R(\mathrm{Inn}(G))$ equals $\mathrm{Aut}(\mathcal{C}_{G}^{+})$. By Theorem 3.3, $\mathrm{Inn}(G)$ is closed in $\mathrm{Aut}(G)$; since $R$ is a topological isomorphism, $R(\mathrm{Inn}(G))$ must in fact be closed, and hence equal to $\mathrm{Aut}(\mathcal{C}_{G}^{+})$. Also $\mathrm{Aut}(\mathcal{C}_{G}^{+})$ is normal in $\mathrm{Aut}(\mathcal{C}_{G})$ because $\mathrm{Inn}(G)$ is normal in $\mathrm{Aut}(G)$. Therefore
$\mathrm{Aut}(G)/\mathrm{Inn}(G)\cong\mathrm{Aut}(\mathcal{C}_{G})/\mathrm{Aut}(\mathcal{C}^{+}_{G}).$
The statement (3) now follows by invoking 2.2. ∎
We recall the following from Definition 1.4. An open subgroup of $G$ is called irreducible if it has no proper open subgroup of finite index; an irreducible subgroup $H$ of $G$ is almost essential if every open subgroup of $G$ contains a finite intersection of conjugates of $H$; $H$ is essential if it is almost essential, and its depth is minimal among the open subgroups of this kind.
We are now ready to prove the results of the introduction on groups with finitely many conjugacy classes of essential subgroups.
Proofs of Theorems 1.5(B) and 1.8(B).
We apply Theorem 5.8. Let $P$ be the property of an open subgroup of being essential. It suffices to show that $P$ is invariant and Borel in the sense of 5.1(a,b). $P$ is clearly invariant. To show that $P$ is Borel, we use the notation from the proof of 5.7, in particular the Borel function $G\to\mathcal{W}(G)$. There is an $L_{\omega_{1},\omega}$-formula $\phi$ with one free variable expressing in $\mathcal{W}(G)$ that a subgroup is essential. Given $U\leqslant_{o}G\leqslant_{c}\mathrm{Sym}(\omega)$ with $G$ oligomorphic, $U$ is essential in $G$ iff $\mathcal{W}(G)\models\phi(\nu_{G}^{-1}(U))$; this condition is Borel. ∎
Borelness of the well studied class of $\mathfrak{G}$-finite groups (Definition 1.6) can be shown in a similar way; we omit the proof.
Proposition 5.9.
Let $G$ is $\mathfrak{G}$-finite. Then $G$ has an essential subgroup.
Proof.
By 4.9 we can assume that $G=\mathrm{Aut}(M)$ and $G$ has a single $1$-orbit. For each $a\in M$, let $H(A)\leqslant G_{a}$ be the smallest open subgroup of $G$ that is a subgroup of $G_{a}$ with finite index. The $H(A)$ is irreducible. Since each open subgroup of $G$ contains the pointwise stabiliser of a finite set, each group of the form $H(A)$ is almost essential in the sense of 1.4. Hence there is an almost essential subgroup of minimal depth, which is then essential by definition. ∎
As indicated in 1.9, we do not know whether a $\mathfrak{G}$-finite oligomorphic group can have infinitely many conjugacy classes of essential subgroups.
Remark 5.10.
As a matter of methodological interest, we can also obtain Theorem 1.5(A) (that the class of oligomorphic groups with no algebraicity is smooth) from the slightly more general form of Theorem 5.8 where the given groups are in a Borel class, making use of 4.6 as before. Let $P$ be the property of an open subgroup that $U=G_{a}$ for some $a\in\omega$. We claim that this property is isomorphism invariant among the groups without algebraicity. For suppose $G,H$ are such groups, and $\Phi\colon G\to H$ is an isomorphism. This corresponds to a bi-interpretation of $M_{G}$ and $M_{H}$ by [1]. The proof of 4.6 shows that this bi-interpretation can be replaced by a bijection $\theta\colon M_{G}\to M_{H}$ preserving definability in both directions, without changing the corresponding topological isomorphism $\Phi$. Clearly $\Phi(G_{a})=H_{\theta(a)}$ for each $a\in\omega$. Thus $P$ is preserved. This also yields that $\mathrm{Aut}(G)$ is quasi-oligomorphic (part of Theorem 1.8(A)), but not that $\mathrm{Out}(G)\cong N_{G}/G$ in 4.7(3).
Note that an open subgroup is irreducible iff it is the pointwise stabiliser of some set $\mathrm{acl}(e)$ where $e\in M^{\mathrm{eq}}$. We provide some model theoretic context to the other condition in the definition of being almost essential in 1.4.
Proposition 5.11.
Suppose that $U$ is an open subgroup of $G=\mathrm{Aut}(M)$. Write $U=G_{e}$, for $e\in M^{\mathrm{eq}}$ and let $S\subseteq M^{\mathrm{eq}}$ be the $G$-orbit containing $e$. Then the following are equivalent.
  1. (i)
    Every open subgroup of $G$ contains a finite intersection of conjugates of $U$;
  2. (ii)
    The permutation group induced by $G$ on $S$ is faithful and closed; that is, the canonical map $r:G\to\mathrm{Sym}(S)$ given by the action of $G$ on $S$ is injective and has closed range;
  3. (iii)
    The structure induced by $M^{\mathrm{eq}}$ on $S$ is bi-interpretable with $M$.
Proof.
(i) $\to$ (ii). Firstly, the kernel of the map $r$ in (ii) is contained in every open subgroup by (i), and so must be trivial. Secondly, for every $a\in M$ there is a finite $B\subseteq S$ such that $a\in\mathrm{dcl}_{M^{\mathrm{eq}}}(B)$. This implies that the group $r(G)$ induced on $S$ is closed: if $r(g_{n})$ ($n\in\omega$) is a convergent sequence in the range of $r$, then the $g_{n}$ converge and the limit $g$ is in $G$; so the $r(g_{n})$ converge to $r(g)$.
(ii) $\to$ (i). $r$ gives a continuous isomorphism between the Polish groups $G$ and $r(G)$, so it is a homeomorphism. Thus, if $H\leqslant G$ is open, then $r(H)$ is open and so contains $G_{(B)}$ for some finite subset $B$ of $S$.
(iii) $\to$ (ii) follows from the above mentioned correspondence of topological isomorphisms and bi-interpretations [1].
(ii) $\to$ (iii). The structure $N$ induced by $M^{\mathrm{eq}}$ on $S$ is bi-definable with the structure obtained by naming all the orbits of the action of $G$ on $S$. So, if the map $r:G\to\mathrm{Sym}(S)$ is injective and has closed range, then $\mathrm{Aut}(N)\cong G$. ∎
5.2. Smoothness for weak elimination of imaginaries
The following is equivalent to the standard definition of weak elimination of imaginaries; see e.g. [9, Lemma 1.3] for a definition as well as a proof of the equivalence.
Definition 5.12.
Let $M$ be $\aleph_{0}$-categorical. We say that $M$ has weak elimination of imaginaries (w.e.i.) if for every open $U\leqslant\mathrm{Aut}(M)$ there is a unique finite algebraically closed set $A\subseteq M$ such that $G_{(A)}\leqslant U\leqslant G_{\{A\}}$.
In the following we reprove the main result of [20, Th 1.1] in our framework (relying on some technical lemmas in [20]), and use this to obtain new information about the groups $\mathrm{Aut}(G)$ where $G=\mathrm{Aut}(M)$ and $M$ has w.e.i.
Corollary 5.13 (to Theorem 5.8).
The class $\mathcal{C}$ of automorphism groups of $\aleph_{0}$-categorical structures that have w.e.i.  is smooth [20]. Furthermore, if $M$ has w.e.i. and $G=\mathrm{Aut}(M)$, then $\mathrm{Aut}(G)$ is quasi-oligomorphic and $\mathrm{Out}(G)$ is profinite.
Proof.
$\mathcal{C}$ is Borel by [20, 2.18]. It suffices to show that $\mathcal{C}$ is a subclass of $\mathcal{B}_{P}$ where $P$ is ‘essential subgroup’. Let $G=\mathrm{Aut}(M)$ where $M$ has w.e.i.
To verify that $G$ has an essential subgroup, by 5.9 it suffices show that $G$ is $\mathfrak{G}$-finite. To see this, suppose a finite set $A\subseteq M$ is algebraically closed. We claim that if $V\subseteq G_{(A)}$ is open and of finite index then $V=G_{(A)}$. We have $V=G_{e}$ for some imaginary $e$. Let $B=\mathrm{acl}(e)\cap M$. Since $M$ has w.e.i. we have $G_{(B)}\leqslant G_{(e)}\leqslant G_{\{B\}}$. So $A\subseteq B$ (as $G_{(B)}\leqslant G_{(A)}$) and $e\in\mathrm{acl}(A)$ (as $G_{(e)}$ is of finite index in $G_{(A)}$). It follows that $A=B$.
The irreducible subgroups of $G$ correspond precisely to the pointwise stabilisers of finite algebraically closed subsets of $M$ by [20, 2.11]. For any fixed depth $k<\omega$ there are only finitely many conjugacy classes of such stabilisers by [20, 2.17]. A fortiori, there are only finitely many conjugacy classes of essential subgroups. ∎
6. Appendix (with D. Evans): an example related to Question 1.9
A more general form of Question 1.9 is the following. Let $G$ be $\mathfrak{G}$-finite and let $k\geqslant 1$. Does $G$ have only finitely many conjugacy classes of almost essential subgroups of depth $k$? This has a negative answer. We thank David Evans for agreeing to include his counterexample in our paper.
We define a theory $T$ in a signature with relation symbols $R_{n}(x;\bar{z},\bar{z}^{\prime}),n<\omega)$ where: $x$ is a single variable, and $\bar{z},\bar{z}^{\prime}$ are tuples of $n$ variables. The theory has axioms saying that for each $x$, $R_{n}(x;\bar{z},\bar{z}^{\prime})$ is either empty, or an equivalence relation on $n$-tuples of distinct elements different from $x$. $T$ has a model companion. A model $M$ of it is simply the Fraisse limit of the finite models of $T$. Then $M$ is a transitive $\aleph_{0}$-categorical structure with trivial algebraic closure; each $R_{n}(a;M,M)$ has infinitely many equivalence classes, that are all infinite.
Now, for each $n$ we have a definable equivalence relation $E_{n}$ on $(n+1)$-tuples $\bar{c},\bar{c}^{\prime}$ of distinct elements of $M$ saying that the first coordinates of $\bar{c},\bar{c}^{\prime}$ are equal (say, to $a$) and the remaining coordinates are $R_{n}(a;M,M)$-related. Any $a\in M$ is in the definable closure of some $E_{n}$-class. So the corresponding sort $S_{n}=M^{(n+1)}/E_{n}$ is such that for any $e\in S_{n}$ we have that $\mathrm{Aut}(M)_{e}$ is almost essential. Furthermore, elements from different sorts $S_{n}$ and $S_{m}$ give non-conjugate almost essential subgroups of the same depth. Notice finally that the $\mathrm{Aut}(M)_{e}$’s have depth $>1$ as they are all properly contained in $\mathrm{Aut}(M)_{a}$, for some $a\in M$.